7. Happy Valley Ice Cream Company starting salary for entry level marketing employees is based on an initial starting salary of the maximized function in Problem No. 4 times $1000:

Initial salary_________________

To entice young employees, they provide a 6% linear commission on sales and all the ice cream you can eat.

Salary______________________

The company’s location is in rural Vermont and to further encourage young employees to move they are providing a $10,000 bonus payment if you keep the job for 1 year. What do you think of this incentive ?

Problem No. 4:

4. Use Linear Programming by Graphing and show a chart with the corners and solution

Maximize f = 2x + 8y

Subject to x + y ≤ 7

2x + y ≤ 12

x + 3y ≤ 15

x ≥0, y≥0
Answer #1

тах f = 2+ + 8 } subject to xty = 7 2х + < 1- xt 39 < Is x , y 7to we will draw these lines on graph ху 2+ $2 13 Y coordinate (X2) Point X coordinate (X1) OR Value of objective function (Z) 0 0 7 two 12 F 6 0 G 4.2 3.6 H05 I fifteenpoints so we have four corner cois).,(3,4),(5,2), (60) - maximum f is at f = 40 (0,5). andG. Hence initial salary is 1000 x 40 = $40,000 on sales 6% linear commission 40000 salary = 40000 t. 6 (0o = 4000 + 2400 = $